OpenStudy (anonymous):

Need help finding the surface area of a hexagonal pyramid. Length: 24 cm,base is 7cm and the sides 24.8 cm

OpenStudy (anonymous):

find the areas of the individual pieces and sum for the hexagon since it's a pyramid i assume its regular so find the area of a triangle in the hexagon then multiply by 6. likewise find area of one of the triangles of the pyramid then multiply by 6.

OpenStudy (anonymous):

So,I'm not sure if I got the answer right but I looked up Surface Area and it's 2b+ph (2 base+ Perimeter times height) My answer came out to be: 228 Does that sound right ?

OpenStudy (dumbcow):

Be careful, that formula only works for finding surface area of prisms. For a pyramid find area of base plus areas of triangle faces. Base is a hexagon which can be tricky, imagine finding area of 6 equilateral triangles of length 7. ->Area of triangle = 1/2 * length*height = 1/2 * 7* (sqrt(3)*7/2) = 49*sqrt(3)/4 ->Area of hexagon base = 6*49*sqrt(3)/4 = 73.5*sqrt(3) Triangle face has length of 7 and height of 24.8 _>Area of triangle = 1/2 * 7 * 24.8=86.8 There are 6 triangle faces so multiply by 6 and then add to base area ->SA = 6*86.8 + 73.5*sqrt(3) = 648

OpenStudy (anonymous):

thanks for helping! I'm having trouble entering your explaination into my calculator. For example: 6* 86.8+73.5*square root =648 did not come out for me.